Alexia bought little bags of nuts to take on her bus trip. She had two bags of cashews, three of pecans and three of pine nuts. If she pulls out two bags at random (without replacement) what is the probability that they will be different kinds of nuts?
Answer by htmentor(1343)

Alexia bought little bags of nuts to take on her bus trip. She had two bags of cashews, three of pecans and three of pine nuts. If she pulls out two bags at random (without replacement) what is the probability that they will be different kinds of nuts?
============================================
The probabilities that they are both the SAME are:
P(cashews) = 2/8*1/7 = 2/56
P(pecans) = 3/8*2/7 = 6/56
P(pine nuts) = 3/8*2/7 = 6/56
The sum of the individual probabilities gives the total probability that they are the same:
P(same) = 2/56 + 6/56 + 6/56 = 14/56
Therefore, the probability they are different = 1 - P(same) = 42/56 = 3/4